the length of a rectangle is twice the width. Find the width if the perimeter is 60 centimeters. Define a variable, and write an equation, and solve.
Answer by checkley77(12844)

L=2W
2L+2W=60
2*2W+2W=60
4W+2W=60
6W=60
W=60/6
W=10 ANS. FOR THE WIDTH.
L=2*10=20 ANS. FOR THE LENGTH.
PROOF:
2*20+2*10=60
40+20=60
60=60